OpenStudy (anonymous):

Find the first partial derivatives of f(x, y) = sin(x − y) at the point (1, 1). A. fx(1, 1) = ??? B. fy(1, 1) = ???

OpenStudy (nikvist):

\[f(x,y)=\sin{(x-y)}\]\[\frac{\partial f}{\partial x}=\cos{(x-y)}\quad,\quad\frac{\partial f}{\partial y}=-\cos{(x-y)}\]\[\frac{\partial f}{\partial x}(1,1)=1\quad,\quad\frac{\partial f}{\partial y}(1,1)=-1\]

OpenStudy (anonymous):

thanks :D